OpenStudy (anonymous):

SIMPLE & EASY QUESTION - MEDAL & FAN what is y(-6/y) -6/y is a fraction

OpenStudy (catlover5925):

if it is easy why are you asking??

OpenStudy (anonymous):

I forget easy stuff about fractions.

OpenStudy (anonymous):

I got -6y/y, not sure if you're suppose to multiple y by everything in the fraction??

OpenStudy (anonymous):

what is ur question?

OpenStudy (anonymous):

What's y *(times) -6/y

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

it's not 6, i'm not that stupid.

OpenStudy (anonymous):

the y cancels the y leaving -6/1 =-6

OpenStudy (anonymous):

x'DDDD. How do they cancel? if it's multiplying.

OpenStudy (anonymous):

y(-6/y) = -6y/y = -6/1 = -6